Primary tabs

Story Archive

Category Title Date Excerpt Attached Images
Current Events Beam delivery update: Development finished; operation resuming 21 October 2014 Today’s scheduled beam development has finished and operation is resuming. The cyclotron is currently delivering ~105 µA to BL1A, 40 µA to BL2A (as the 2A current is still being stepped up), and ~95 µA to BL2C4.
Current Events Beam delivery update: Beam development tomorrow 20 October 2014 The cyclotron is currently delivering ~110 µA to BL1A and 50 µA to BL2A. This is expected to continue until the start of tomorrow’s planned beam development shift. Development is scheduled to run from 08:30 until roughly 20:00, at which time normal operation will be restored.
Current Events Beam delivery update: major power bump Saturday 19 October 2014 As many of you already know, the lab experienced a major power bump at about 09:45 yesterday. This had a big impact on almost every accelerator-related system on site: vacuum, cryogenics, RF, controls, cooling systems, building services, etc.
Current Events Beam delivery update 17 October 2014 The cyclotron is presently delivering ~110 µA to BL1A and 50 µA to BL2A; BL2C1 is set up for proton therapy tomorrow morning. High-current delivery to BL1A and BL2A, interrupted each morning for proton therapy, is expected to continue through the weekend.
Current Events October 17 ARIEL Snapshot 17 October 2014 Machine Protection System tests will start in ~2 weeks. Beam dump lower centre block to be sent out for milling. Interlock Specification document under final review.
Current Events Canada's Big Science Leaders Converge 17 October 2014 Leaders from four of Canada’s big science facilities joined together at Canadian Science Policy Conference in Halifax today to discuss how large research infrastructure provides significant societal and economic benefit to Canada.
Current Events Beam delivery update: Proton therapy starting tomorrow 16 October 2014 The cyclotron is delivering ~110 µA to BL1A and 50 µA to BL2A. This is expected to continue until the start of proton therapy at 08:00 tomorrow morning. BL2C1 was set up for PT this afternoon.
Current Events Beam delivery update 15 October 2014 The cyclotron is now running pretty much normally, delivering ~110 µA to BL1A and 50 µA to BL2A. The STF group hasn’t yet decided if they’re going to request beam overnight or if BL2C will stay off until tomorrow's switchover to 2C1.
Current Events Award Recognizes Italian-Canadian Ties 15 October 2014 In a private ceremony yesterday, Dr. Emanuele Fiore, Scientific Attaché of the Embassy of Italy, was presented with TRIUMF’s inaugural Patron of Science award. The award recognizes selfless and meritorious contributions of an individual outside the organization.
Current Events Further to last night's post… 14 October 2014 The main magnet is on, the RF has been tested, and inflector/deflector conditioning is underway. The present hope is to inject beam some time later today for full production tomorrow as originally scheduled.

Pages

Subscribe to Recent Stories